I need help translating and solving by system of linear equation please help.
In an isosceles triangle, each base angle is 10degrees less than two times the third angle. find the measures of all three angles.

Answer by mananth(16946)

third angle =x
base angle = 2x-10
2x-10 +2x-10+x=180 ( sum of angles of trianle)
5x-20=180
5x=200
x=40 deg
2x-10 = 2*40-10=>70
base angle = 70
check
70+70+40=180
m.ananth@hotmail.ca